1. Propane and oxygen react according to the following equation.

C3H8(g) + 5O2(g) 🡪 3CO2(g) + 4H2O(g)

Calculate the volume of carbon dioxide and water vapour produced and the volume of oxygen remaining, when 20.0 L of propane reacts with 120.0 L of oxygen at STP.

  1.  What volume of carbon dioxide will be produced in the reaction of 100.0mL of 1.1M sodium carbonate with 250.0mL 1.15M hydrochloric acid if its yield is 65%?      

 Na2CO3(aq)  +  2HCl(aq)   🡪   2NaCl(aq)   +    H2O(l)   +   CO2(g)



1
Expert's answer
2021-07-30T14:07:01-0400

The volumes are all measured at the same conditions of temperature and pressure. This implies that mole ratios may be interpreted as volume ratios. Have


Volume of "Co_2 ={ 3}\u00d7 {20}=60" L


Volume of "H_2O=\frac{120×4}{5}=" 96L



"Na_2CO_3(aq) + 2HCl(aq) ---> 2NaCl(aq) + H_2O(l) + CO_2(g)"


Moles of sodium carbonate ="\frac{100×1.1}{1000}=0.11moles"


Moles of HCL"=\frac{250×1.15}{1000}=0.2875moles"


Moles of "CO_2=0.11"